### v1

# Explainable Compliance Detection with Multi-Hop Natural Language Inference on Assurance Case Structure

Abstract
Ensuring complex systems meet regulations typically requires checking the validity of assurance cases through a claim-argument-evidence framework. Some challenges in this process include the complicated nature of legal and technical texts, the need for model explanations, and limited access to assurance case data. We propose a compliance detection approach based on Natural Language Inference (NLI): EXplainable CompLiance detection with Argumentative Inference of Multi-hop reasoning (EXCLAIM). We formulate the claim-argument-evidence structure of an assurance case as a multi-hop inference for explainable and traceable compliance detection. We address the limited number of assurance cases by generating them using large language models (LLMs). We introduce metrics that measure the coverage and structural consistency. We demonstrate the effectiveness of the generated assurance case from GDPR requirements in a multi-hop inference task as a case study. Our results highlight the potential of NLI-based approaches in automating the regulatory compliance process.
Explainable Compliance Detection with Multi-Hop Natural Language Inference on Assurance Case Structure
Fariz Ikhwantri Simula Research Laboratory Oslo, Norway fariz@simula.no                        Dusica Marijan Simula Research Laboratory Oslo, Norway dusica@simula.no
1 Introduction
Compliance detection is a task to ensure that complex systems such as software adhere to established standards, regulations, and best practices (Saeidi et al., 2021; Castellanos Ardila et al., 2022). Compliance detection aims to identify potential violations early in the development lifecycle, reducing legal risks and costly rework (Duvall et al., 2007).
Figure 1: Claim-Argument-Evidence structure of an assurance case can be represented as NLI with direct and multi-hop (indirect) connections
However, ensuring compliance is challenging due to the complexity of legal documents. Navigating privacy policy documents requires two sets of unique expertise in legal and technical domains. This process can be tedious for a legal expert, an auditor, or an engineer. This motivates many experts to automate the process of compliance detection. However, automating this process also means the AI system needs to be transparent. This demand makes it challenging to adapt to the black-box AI system.
Previous study (Cejas et al., 2023) uses inherently interpretable rule-based graph matching, but it still struggles with linguistic variability, ambiguous legal phrasing, and scalability to new regulations. A recent study (Azeem and Abualhaija, 2024) explores statistical machine learning and language model methods for the compliance detection task as a text classification problem. However, it does not consider transparency and explainability factors such as model interpretability and reasoning.
To address the transparency and robustness factors of detecting risk and safety, experts use assurance cases (Rhodes et al., 2010; Mansourov and Campara, 2011) as a structured argument approach to audit complex software products, such as those in aircraft or autonomous vehicles, against standard requirements. This is to ensure compliance with legal and technical standards. Assurance case is a systematic collection of arguments and supporting evidence to demonstrate a certain level of confidence that a product meets particular claims related to its requirements (Netkachova et al., 2015; Piovesan and Griffor, 2017). These assurance cases can be created and updated throughout the development lifecycle (Ross et al., 2022).
Figure 2: The framework of our explainable compliance detection with multi-hop NLI on Claim-Argument-Evidence (CAE) structure of an Assurance Case.
Despite its significance, obtaining and using assurance case data presents important challenges. One of the foremost challenges is the lack of publicly available datasets. Another challenge is the high cost of acquiring and annotating assurance case data. Developing and validating assurance cases requires significant domain expertise, making large-scale annotation expensive and time-consuming.
Recently Odu et al. (2025) proposed to instantiate the Assurance Case patterns of Goal Structuring Notation (GSN) (Kelly and Weaver, 2004) with Large Language Models (LLMs). They assess the data produced by LLMs by utilising shallow lexical metrics from the text generation task. Chen et al. (2025) used LLMs to create Assurance Cases and evaluated them using constraint solvers from Formal Methods (Jaffar and Maher, 1994). However, constraint solvers cannot handle every assurance case completely, as they frequently involve qualitative reasoning, contextual relationships, and uncertainties that extend beyond strict logical or mathematical boundaries.
One possible approach is to use deductive reasoning (Rushby et al., 2015) on the assurance case. Deductive reasoning in the assurance case derives conclusions from structured claims and arguments from evidence (Bloomfield and Bishop, 2009). This reasoning can be framed in the Natural Language Inference (NLI) task (Dagan et al., 2005; Bowman et al., 2015) to determine if a hypothesis follows from a premise. If there is a set of premises, multi-hop inference (Jansen, 2018) is needed for reasoning over interconnected claims and evidence. This makes multi-hop inference suitable for assurance cases that link high-level system properties to low-level evidence, like test results.
This paper proposes EXplainable CompLiance detection with Argumentative Inference of Multi-hop reasoning (EXCLAIM) based on Assurance Case. Figure 2 illustrates our workflow that explains this paper’s contribution, which provides a summary of this paper. From the figure, we use pre-trained transformer (Devlin et al., 2019) models to learn premise-hypothesis pairs for deductive multi-hop reasoning to perform compliance detection.
Specifically, our contributions address challenges in (i) explainable compliance detection and (ii) the lack of assurance case data and advanced reasoning, as follows:
  • We propose compliance detection based on Natural Language Inference. We formulate the claim-argument-evidence as a discourse structure which can be used to develop a multi-hop inference model for explainable tracing of the compliance detection model’s output (Figure 1).
  • We propose evaluation methods to measure the assurance case generated by large language models (LLMs) to validate the reliability of instance-level consistency and structural coherence. We conduct a study by generating Claim Argument Evidence Assurance Case data with open source and proprietary LLMs from GDPR requirements as the context prompt from a previous study (Cejas et al., 2023). We analyse the generated Assurance Case data by using our proposed metrics.
2 Related Work
Compliance detection in Requirements
Compliance with legal regulations and industry standards has become necessary as software systems, such as finance and healthcare, become more integrated into people’s daily lives. To ensure compliance, software requirements must be verified to align with applicable laws, regulations, standards, and organisational policies, for example, in data protection laws (e.g., GDPR, CCPA), cybersecurity (e.g., IEC 62443, NIST 800-53), and safety-critical standards (e.g., HIPAA). This process is known as compliance detection.
Recent studies have extensively evaluated statistical machine learning and Language model approaches as text classification tasks for compliance detection (Azeem and Abualhaija, 2024) on GDPR. Their study reveals that text classification methods outperform traditional rule-based graph matching (Cejas et al., 2023), offering greater flexibility and adaptability in handling various types of data. However, they do not address crucial aspects of explainability, i.e. the reasoning processes behind the model predictions. This lack of focus on explainability raises concerns about the transparency and trustworthiness of the employed automated system.
Explainable AI
Explanation of model prediction provides the rationale behind a model’s decision. The rationale can be obtained by analysing the contribution of individual features to the decision-making process (Molnar, 2025). Model explanations can be either inherent or post-hoc. Shallow models like linear regression are inherently interpretable, whereas deep learning and LLMs typically require post-hoc methods such as gradient-based (Simonyan et al., 2014; Sundararajan et al., 2017) or input-perturbation approaches (Ribeiro et al., 2016; Lundberg and Lee, 2017). In addition, the explanation needs to be evaluated by plausibility (Hollenstein and Beinborn, 2021; Eberle et al., 2022; Ikhwantri et al., 2023, 2024) or faithfulness (Atanasova et al., 2020; Bastings et al., 2022).
3 Methodologies
This section describes our proposed methodologies for forming compliance detection as a natural language inference task and extending it to multi-hop reasoning from the assurance case structure.
3.1 Compliance Detection as Natural Language Inference
Req. Premise (Req) Hypothesis (DPA) DPA-ID Label
R18 The processor shall assist the controller in consulting the supervisory authorities prior to processing … to mitigate the risk. (Art. 28(3)(f), Art. 36) At controller’s request and at the controller’s reasonable expense on a time and materials basis, … fulfilling any … Applicable Data Protection Law. Online 100 entailment
Table 1: Compliance Detection as NLI task where GDPR requirement is the premise and Data Processing Agreement is the hypothesis
Natural Language Inference (NLI) is a core task in natural language understanding that involves determining the logical relationship between two textual statements (Dagan et al., 2005). We propose compliance detection by devising the requirements as premises and evidence as hypotheses. This method allows the model to learn from the verbalisation of requirements text instead of discrete classes compared to previous studies, as text classification (Azeem and Abualhaija, 2024).
For example, in Table 1, the inputs are GDPR requirements, which served as the premise, and the Data Processing Agreement (DPA) served as the hypothesis. The entailment label means requirements and DPA belong to the same GDPR requirement class and comply, while the non-entailment label means requirements and DPA do not belong to the same GDPR requirement class, i.e. neutral or contradict.
This is similar to requirements mapping (Fazelnia et al., 2024) for requirements classification approaches, which are based on few-shot (Wang et al., 2021) or zero-shot scenario (Sainz et al., 2021). The output is similar to binary classification approaches (Azeem and Abualhaija, 2024). However, NLI only requires a model instead of $n$ models for binary classification, where $n$ is the number of requirements.
3.2 Assurance Case Structure as Multi-hop Inference
In the assurance case, the claim-argument-evidence (CAE) structure could be leveraged to construct premise-hypothesis text pairs in the NLI task. These CAE trees provide a formulation for deductive inference, allowing for sequential reasoning, where multiple premises collectively support a conclusion.
We construct multi-hop premise-hypothesis pairs where a hypothesis is inferred from the indirect connection of a sub-claim to evidence or the main claim to an argument sub-claim. An example can be seen in Figure 1. From the figure, the parent node in the assurance case structure serves as the premise. The child node, which builds upon or supports the parent, becomes the hypothesis. These structured relationships are extracted from assurance cases and transformed into a format suitable for training NLI models. This approach enables our systems to perform more complex inferences in a simple NLI model as a strong baseline.
4 Dataset
4.1 GDPR-DPA
We use the compliance detection dataset from the previous study (Azeem and Abualhaija, 2024) for the NLI task on the real-world human-annotated dataset  [^1] https://zenodo.org/records/11047441. We called this dataset as GDPR-DPA from General Data Protection Regulation-Data Processing Agreement (GDPR-DPA). The details of the dataset statistics can be seen in Table 2. The number of instances in training set is larger than original data due to pairing of negative samples between non-matching premise (requirements) and the hypothesis (DPA). We set the sampling rate 0.1 to add the negative pairs within iteration of unique hypotheses times unique premises.
Statistics Train Test
Instances 362317 1511
Unique ID 61 8
Unique premise 45 45
Unique hypothesis 9820 1364
Unique target 45 17
Table 2: Dataset Statistics for NLI-Based Compliance Detection on GDPR-DPA Dataset: Train and Test
4.2 LLM Generated Data
We use the GDPR requirements compiled by Cejas et al. (2023) for the prompt to generate assurance cases. There are 45 requirements, but only 20 are considered for compliance detection (Azeem and Abualhaija, 2024). We generated an Assurance case using three proprietary LLMs, namely gpt3.5-Turbo-4k, gpt4o-8k and gpt-4o. From open source LLMs, we use models such as Llama3 (3.2-1B, 3.1-8B, 3.1-70B) (Grattafiori et al., 2024), Qwen2.5 (7B, 14B, 72B) (Yang et al., 2024), Gemma (7B, 9B, 27B) (Team et al., 2024), and phi4 (Abdin et al., 2024).
4.3 Evaluating LLMs generated Assurance Case
A recent study Odu et al. (2025) used Goal-Structured Notation to generate assurance cases from five products. This structure differs from the Claim-Argument-Evidence structure. In addition, their methods rely on the availability of gold assurance, which is not always available and is hard to obtain. Their work evaluates the LLMs’ generated data using shallow lexical metrics such as the BLEU score from the machine translation task. This lexical metric does not give insight into the higher-level view of the generated data.
Flat metric
We proposed Flat metrics, which assess the assurance case at a high level without analysing textual content or internal relationships between elements. This approach provides a quantitative assessment based on the structural components of the generated assurance case.
We measure the quantitative assessment by calculating the absolute difference for each element type, such as the main claim, sub-claim, argument claim, and evidence from two assurance cases of the same requirements. Flat metrics provide an initial assessment of coverage, balance of elements, and consistency. Concretely, we can detect possible redundancy across assurance cases because they focus on quantitative elements without regard for the structures.
Structured metric
Assurance cases are structured arguments, often represented as graphs or trees. We need to consider the graph’s structure to assess the quality of assurance cases generated by LLMs. This means analysing the hierarchical or graph-based relationships between claims, arguments, and evidence in an assurance case.
One possible approach to measuring the similarity between two graphs is Graph Edit Distance (GED). GED is a measure of dissimilarity between two graphs. This metric quantifies the minimum number of edit operations required to transform one graph into another. A lower value indicates higher similarity, while a higher value indicates lower similarity. The value of zero indicates the two graphs are isomorphic.
Data Analysis
We divide the Flat and Structured metric analysis into intra-model and inter-model comparison to evaluate the claim-argument-evidence (CAE) structure generated by LLMs.
Intra-model comparison evaluates how consistently a single model performs across multiple inferences, similar to intra-annotator agreement in human data labelling. On the other hand, inter-model comparison examines the agreement between the outputs of different models.
Table 3 shows the difference between four types of instances generated by LLMs. In flat metric, generated instances of type MainClaim, SubClaim and ArgumentClaim at the instance levels have lower differences, which means high consistency. Overall, models tend to diverge in evidence instances. This means that efforts can be made to check the validity of the assurance case structure by checking the coverage of evidence rather than the argument and SubClaim.
In structured metrics, Gemma2 models(v2-9b-it, v2-27b-it) and phi4 generate a consistent structure, which is shown by their intra-model structured metrics (GED) results. The lowest inter-model value is a 3.26 GED score between the phi4 and chatGPT4o models.
Type Intra-model Inter-model
MainClaim 0.00 0.00
SubClaim 0.08 0.24
ArgumentClaim 0.16 0.32
ArgumentSubClaim 0.37 0.90
Evidence 0.89 2.90
Table 3: Flat-metrics: Absolute Count Difference for Intra- and Inter-Model Agreement
Figure 3: Graph Edit Distance between intra and inter-model generated assurance case structure.
Table 4 describe the dataset statistics details of final LLMs generated assurance cased as multi-hop inference task.
Train Test
All 103838 12875
#1_hop 12241 3150
#2_hop 10725 2723
#3_hop 9136 2286
#4_hop 6192 1495
Table 4: Dataset statistics of Generated Assurance Case by LLMs in Evaluating LLM Generated Data
5 Experiments Setup
In this section, we explain the experiment setup for evaluating compliance detection as a natural language inference task on the GDPR-DPA dataset and the multi-hop inference from the assurance case data generated by LLMs.
5.1 Research Questions
In this paper, we aim to answer the following questions:
RQ1
How does compliance detection as NLI performance compare to the previous approach?
RQ2
How does the model’s prediction on compliance detection output affect its explanation?
RQ3
What is the model’s performance in compliance detection on the unseen requirements?
RQ4
Does the synthetic reasoning improve the model’s faithfulness in detecting compliance with the unseen requirements?
5.2 Faithfulness Metrics
Faithfulness refers to the degree to which the model relies on the output of local explanation (rationales) to make its prediction. It ensure the provided rationales are not just plausible but also truthful representations of the model’s inner workings (Jacovi and Goldberg, 2020). There are two faithfulness metrics we consider in this work.
Comprehensiveness
evaluates whether all necessary rationales for making a prediction have been chosen (DeYoung et al., 2020; Atanasova et al., 2020). This can be done by creating a contrasting example for x, denoted as $\widetilde{x}$, which is x without the predicted rationales r. In a classification context, let $m{(x)}_{j}$ represent the original model prediction for the output class $j$. Next, we examine the predicted probability from the model for the same class after eliminating the supporting rationales. The model should demonstrate reduced confidence in its prediction once the rationales are removed from x. This can be expressed as $\mathrm{Compr}=m\left(x\right)_{j}-m\left(x|r\right)_{j}$, where ${(x|r)}_{j}$ is the sentence x where tokens in r are removed. This metric indicates that a greater difference between the original sentence and the modified version, excluding the specified tokens, is preferable. A higher value signifies more significant alterations and, consequently, a more effective reduction of the input while retaining essential meaning.
Sufficiency
evaluates whether the identified rationales are adequate for the model to make accurate predictions (DeYoung et al., 2020; Atanasova et al., 2020). It measures the model’s confidence in its prediction when only the rationales are retained, compared to the original input. The sufficiency metric is defined as $\mathrm{Suff}=m\left(x\right)_{j}-m\left(r\right)_{j}$, where $(m\left(x\right)_{j})$ represents the model’s original prediction confidence for the output class $(\ j\ )$, and $(m\left(r\right)_{j})$ represents the model’s prediction confidence when only the rationales $(\ r\ )$ are retained. In this metric, a value close to zero is better, indicating that the model can still make confident predictions with just the rationales, demonstrating their Sufficiency.
6 Experiment Scenarios
6.1 NLI-Based Compliance Detection
To address the challenges of legal text ambiguity, domain specificity, and evolving regulations in compliance detection, we explore transfer learning methods and compare the results with a previous study (Azeem and Abualhaija, 2024) to answer [RQ1]. In this scenario, we compare three different transfer learning algorithms, namely Fine-tuning (FT), Linear-Probing-Fine-tuning (LP-FT) (Kumar et al., 2022) and Low-Rank Adaptation (LoRA) (Hu et al., 2022).
Our approach considers two main pre-trained language models (LM) categories: general-purpose and domain-specific legal models. For general domain, we consider BERT-large (Devlin et al., 2019) (BERT), RoBERTa-large (Liu et al., 2019) (RoBERTa). For domain-specific legal models, we consider Legal-BERT-base (Chalkidis et al., 2020) (Legal-BERT) as the counterpart of BERT and Legal-RoBERTa (Chalkidis et al., 2023) as the counterpart of RoBERTa. In addition, we also explore the decoder-based LM, such as GPT-2-XL222openai-community/gpt2-xl and Llama-3.2-1B (Grattafiori et al., 2024), to classify the entailment as a text generation baseline.
To answer [RQ2], we explore how model faithfulness influences compliance detection output. We evaluate four post-hoc explanation methods, such as Gradient (Grad) (Simonyan et al., 2014), Integrated gradient (IG)(Sundararajan et al., 2017), LIME (Ribeiro et al., 2016) and Partition SHAP (Part SHAP) (Lundberg and Lee, 2017). In this paper, we use the ferret library (Attanasio et al., 2023) implementation to measure the model’s faithfulness. The faithfulness metric output is a score of the area over the perturbation curve (AOPC). AOPC is obtained by aggregating the mean of the Comprehensiveness or sufficiency score over bins of 10% which are applied to the models.
6.2 Multi-Hop Inference
We evaluate whether transformer-based language models can perform deductive inference on CAE tree-based assurance cases. We explicitly append indirect premises to the model input as a simple baseline. Appending indirect premises allows the models to perform multi-hop inference. We aim to assess whether explicit intermediate reasoning steps improves inference performance compared to a model that only sees the indirect premise and hypothesis.
In this scenario, we split the train and test of the generated assurance case based on the requirement categories to answer [RQ3]. From the 19 requirements, we used 15 unique requirements for training data and requirements for test data. We use the same models but decided to focus on only fine-tuning to focus on analysing the explanation improvement to answer [RQ4].
7 Results and Discussion
Model Method pre rec f1 f2
Azeem and Abualhaija (2024)
BERT* Binary-FT .82 .82 .82 .82
Multi-FT .83 .85 .84 .84
RoBERTa* Binary-FT .84 .63 .74 .66
Multi-FT .81 .90 .85 .87
This Work
FT .81 .85 .83 .83
BERT LP-FT .91 .76 .81 .79
LoRA .81 .82 .81 .82
FT .90 .83 .86 .84
Legal-BERT LP-FT .89 .75 .80 .77
LoRA .75 .56 .58 .59
FT .89 .72 .78 .75
RoBERTA LP-FT .50 .50 .50 .50
LoRA .80 .86 .82 .85
FT .50 .50 .50 .50
Legal-RoBERTA LP-FT .50 .50 .50 .50
LoRA .91 .86 .88 .87
GPT-2-XL (1.5B) Zero-Shot .49 .48 .48 .48
One-Shot .53 .58 .55 .57
Llama-3.2 (1B) Zero-Shot .49 .49 .49 .49
One-Shot .52 .53 .52 .53
Table 5: Pre-trained language model trained on Compliance detection as NLI task. Previous work use base model size (108M) vs ours mostly use large (334M) except for Legal-BERT
NLI-Based Compliance Detection
Table 5 shows compliance detection as NLI performance compared to previous approaches. We report the precision (pre), recall (rec), and f2-score. We follow previous approaches for a fair comparison to report the f2-score because it puts more weight on recall than precision. The BERT-FT approach performs better than the BERT with binary classification fine-tuning (Binary FT). BERT with multi-class fine-tuning (Multi-FT) performs better than our best results. However, we do not find the decoder-based LM (GPT-2-XL and Llama-3.2-1B) performance satisfactory.
The recall performance between our BERT-FT and the previous approach, BERT with multi-class classification fine-tuning (Multi-FT), is competitive. Our approach is better than BERT binary classification fine-tuning (Binary-FT). The answer to RQ1 "How does compliance detection as NLI performance compare to the previous approach?" is that our best performing approach performs comparably to previous approaches (Binary-FT). However, our approach can provide better explainability and traceability to verbalisation requirements.
Table 6 shows comprehensiveness and sufficiency scores for each model and explainer combination. Overall, LIME methods perform best in terms of the comprehensiveness metric. Regarding the sufficiency metric, Legal-BERT-FT with Part-SHAP shows closer values to zero.
AOPC Comprehensiveness AOPC Sufficiency
Model Explainer Overall Correct Incorrect Overall Correct Incorrect
BERT-FT IG .276(.26) .606(.17)* .168(.18) -.009(.19) .188(.19)* -.073(.13)
Grad (x Input) .215(.26) .467(.26)* .133(.20) .047(.23) .298(.24)* -.035(.15)
LIME .414(.25) .694(.13)* .323(.21) -.058(.15) .048(.15)* -.093(.13)
PartSHAP .367(.25) .671(.13)* .267(.20) -.059(.14) .038(.14)* -.091(.13)
Legal-BERT-FT IG .137(.30) .760(.14)* .021(.13) .032(.24) .518(.17)* -.059(.10)
Grad (x Input) .072(.27) .658(.16)* -.038(.09) .069(.31) .752(.14)* -.058(.10)
LIME .210(.30) .764(.15)* .107(.19) .018(.22) .445(.17)* -.062(.10)
PartSHAP .243(.30) .759(.15)* .147(.21) .006(.19) .367(.19)* -.061(.10)
RoBERTa-LoRA IG .450(.36) .666(.15)* -.044(.15) .376(.38) .599(.18)* -.135(.17)
Grad (x Input) .314(.34) .509(.19)* -.133(.15) .474(.42) .729(.16)* -.112(.16)
LIME .629(.31) .807(.12)* .218(.17) .214(.30) .374(.19)* -.154(.16)
PartSHAP .567(.36) .784(.13)* .066(.16) .237(.31) .412(.17)* -.166(.16)
Legal-RoBERTa-LoRA IG .446(.30) .587(.17)* -.012(.14) .392(.35) .562(.17)* -.167(.16)
Grad (x Input) .298(.31) .423(.23)* -.107(.14) .515(.41) .720(.18)* -.155(.15)
LIME .672(.26) .791(.13)* .283(.21) .154(.25) .258(.17)* -.187(.15)
PartSHAP .612(.27) .742(.12)* .190(.20) .177(.25) .287(.15)* -.183(.15)
[…]
Further analysis shows that correct model predictions generally have higher comprehensiveness scores than incorrect ones, suggesting that faithfulness is associated with predictive accuracy. While the sufficiency metric’s on the correct scores should be closer to zero than the incorrect ones, we argue that it might not be a reliable indicator of completeness in legal compliance checking. Relying only on salient words can lead to ambiguous and misaligned hypotheses with real-world legal reasoning. However, the sufficiency metric might be more relevant to higher-level representation, such as when sentences or paragraphs are kept within documents instead of tokens.
Interestingly, we do not find any significant effect of in-domain pre-training on model faithfulness, indicating that domain-specific training may improve accuracy but does not necessarily enhance interpretability or explanation quality.
The answer to our RQ2 "How does the model’s prediction on compliance detection output affect its explanation?" is that the model’s correct prediction affects explanation on Comprehensiveness. In contrast, the model’s incorrect prediction relates more to the sufficiency metric than the correct prediction.
Multi-Hop Inference
Figure 4 shows results from our experiment in multi-hop inference from SACs. From the figure, we can observe the accuracy drop as the number of hops increases for the indirect model (wo_chain). This result suggests that deep reasoning structures are challenging for NLI models without explicit intermediate steps. The answer to RQ3 "What is the model’s performance in compliance detection on the unseen requirements?" is that the explicit model’s performance is better on unseen requirements than the implicit model’s.
12340.70.70.70.70.750.750.750.750.80.80.80.80.850.850.850.850.90.90.90.90.950.950.950.951111HopMacro F1 ScoreBERT-large, wo_chainBERT-large, chainRoBERTa-large, wo_chainRoBERTa-large, chainLegal-BERT, wo_chainLegal-BERT, chainLegal-RoBERTa, wo_chainLegal-RoBERTa, chain
Figure 4: F1 Scores for Different Models Across Various Hop Difficulties
We analyse the post-hoc explanation of models with (chain) vs without chain (wo_chain). Figure 5 shows the faithfulness metrics difference between without chain and with chain. In the Comprehensiveness metric, ideally, the score should increase from the model without the chain to the model with the chain. These results were observed in the Legal-BERT model and RoBERTa. There are some exception however with the BERT model and RoBERTa models with PartSHAP. In the Sufficiency metric, ideally, the score should decrease from the model without the chain to the model with the chain. These results were only observed in the RoBERTa model with LIME and a slight decrease in the RoBERTa model with IG. The answer to RQ4 "Does the synthetic reasoning improve the model’s faithfulness in detecting compliance with the unseen requirements?" is that the synthetic reasoning improves the model’s Comprehensiveness (by increasing value) in some models and explainer combinations. In the sufficiency metric, the synthetic reasoning not improves (by decreasing value) the explicit chain models, except for the RoBERTa model with LIME and IG.
Figure 5: Slope Chart of AOPC Compr by Model and Method
8 Conclusion
This work introduces a method for compliance detection using multi-hop inferences based on generated assurance cases using LLMs. We decompose long text requirements into claims, arguments, and evidence (CAE) of assurance cases. To assess the effectiveness of our approach, we first compared it against prior methods (RQ1). Formulating compliance detection as the NLI task yields competitive performance and better recall in most case. We then examined the relationship between model predictions and their explanations (RQ2). We observed that the model’s correct prediction affects the explanation based on the Comprehensiveness metric. In contrast, the model’s incorrect prediction relates more to the Sufficiency metric
In evaluating generalisation (RQ3), our explicit reasoning methods demonstrated superior performance on unseen requirements, highlighting the benefit of synthetic data from LLMs. Finally, we explored whether synthetic reasoning steps improve model faithfulness (RQ4). By comparing explicit vs implicit models, we found that incorporating intermediate reasoning chains increases model explainability based on the faithfulness metrics. Our results suggest that generating assurance cases with LLMs offers a promising direction for robust and explainable compliance detection.
9 Limitations
Our study on compliance detection using Natural Language Inference (NLI) and multi-hop reasoning has several potential limitations that might be considered.
Our evaluation relies on assurance cases generated by LLMs, which may introduce hallucinations, inconsistencies, or biases. While we propose evaluation metrics to measure structural coherence and instance-level consistency, these do not fully ensure logical correctness.
Our second experiment uses synthetic assurance case data due to the limited availability of expert-annotated real-world cases. This may affect the generalizability of our results when applied to actual regulatory audits and compliance verification
While focusing on GDPR as a use case, our approach may not generalise seamlessly to other compliance domains such as HIPAA, ISO 27001, or SOX. Other regulatory texts may require domain-specific hyper-parameters, different architecture or transfer learning methods.
The flat and structured metrics we propose measure structural coherence and textual consistency generated by LLMs. They do not aim to fully assess logical validity, factual correctness, or legal soundness. Future work can integrate our evaluation with formal verification or uncertainty quantification techniques to enhance reliability.
10 Ethical Considerations
Our second study experiments on synthetic assurance case data generated by large language models (LLMs) due to limited real-world datasets. While necessary for scalability, this introduces risks of hallucinated content, biased reasoning, or oversimplified logic that may not reflect actual regulatory processes.
The proposed system is not intended to replace legal or expert judgment, and we caution against its use in high-stakes compliance decisions without human oversight. Although we use GDPR as a case study, generalising to other regulatory domains (e.g., HIPAA or SOX) may require domain-specific adjustments.
We encourage responsible use of our methods in contexts where outputs are critically evaluated, and stress the importance of integrating expert validation and, where possible, formal verification.
Appendix A Appendix
A.1 NLI-Based Compliance Detection on GDPR-DPA Dataset experiments
FT
All models are fine-tuned for six epochs, 32 batches, 1e-05 learning rate, and weight decay 0.01. The models are optimised with a 1e-05 learning rate, and the weights decay by 0.01.
LP-FT
All models are fine-tuned for three epochs in the probing phase and three in the fine-tuning phase. The batch size is 32. The models are optimised with a 1e-05 learning rate, and the weights decay by 0.01.
LoRA
All models are fine-tuned for six epochs, 32 batches, 1e-05 learning rate, and weight decay 0.01. The models are optimised with a 1e-05 learning rate, and the weights decay by 0.01. In LoRA parameters, we target these named parameters: ["attention.self.query", "attention.self.key", "attention.self.value", "attention.output.dense", "intermediate.dense"]
Zero-shot & One-shot
For Decoder-based LMs, we used zero-shot and one-shot prompts to predict the entailment. The prompt template for zero-shot is:
Zero-shot Prompt "Below is a Natural Language Inference (NLI) task for compliance detection in general data privacy regulation domain."
"give answer in either ’entailment’ or ’not entailment’"
f"Premise: {premise}"
f"Hypothesis: {hypothesis}"
"Answer:"
The prompt template for one-shot is:
One-shot Prompt "Below is a Natural Language Inference (NLI) task for compliance detection in general data privacy regulation domain.
" f "Example 1:
Premise: {pos_example[0]}"
f"Hypothesis: {pos_example[1]}"
"Answer: entailment"
f "Example 2:
Premise: {neg_example[0]}"
f"Hypothesis: {neg_example[1]}"
"Answer: not entailment"
f"Premise: {premise}"
f"Hypothesis: {hypothesis}"
"Answer:"
Post-hoc Explanation Analysis
Figure 6: Comprehensiveness and Sufficiency Metrics by Model and Explainer
(a) Score distribution of Correct vs Incorrect for BERT
(b) Score distribution of Correct vs Incorrect for Legal-BERT
Figure 7: Score distributions of Correct vs Incorrect for BERT and Legal-BERT
(a) Score distribution of Correct vs Incorrect of RoBERTa-LoRA faithfulness metrics
(b) Score distribution of Correct vs Incorrect of Legal-RoBERTa-LoRA faithfulness metrics
Figure 8: Score distributions of Correct vs Incorrect for RoBERTa and Legal-RoBERTa
A.2 LLM Generated Data
LLM Prompt
In this paper, we use the following system and user role prompt:
LLM System Message You are a legal expert in privacy and security issues. Your duty is to produce a thorough Data Processing Agreement Assurance Case from General Data Protection Regulation legal requirements. {assurance_case_definition}
LLM Prompt The requirement is {requirement_id}: {requirement_name} Requirement Description {requirement_description} Rationale and Supplemental Guidance {requirement_rationale} Give the output in {format}.
Generated text to Assurance Case in JSON format Success rate
Table 7: LLMs generation Error Summary
Metric Count
Number of error files 587
Total files 950
Table 8: Models with Error Output and Success Percentage (Total = 5 calls $\times$ 20 requirements = 100 cases)
Model Err Cnt Success %
Qwen2.5-72B-Instruct 84 16%
Llama-3.1-8B-Instruct 92 8%
gemma-2-27b-it 95 5%
Llama-3.1-70B-Instruct 91 9%
phi-4 5 95%
Qwen2.5-14B-Instruc t 32 68%
Llama-3.2-3B-Instruct 8 92%
Qwen2.5-7B-Instruct 0 100%
gemma-7b-it 95 5%
gemma-2-9b-it 85 15%
gpt35-turbo-4k 4 4%
gpt4o-8k 6 6%
gpt4o 2 2%
We repeat five calls each in the assurance case generation process by 13 LLMs with 20 requirements. In total, there are 1,300 instances of JSON files. However, only around 25% of files are valid JSON files in the first attempt. We make another call to the ChatGPT4o for each malformed JSON, which has a 100% success rate in formatting JSON. Finally, we have 1000 valid Assurance Case JSON files, where 727 files are from Open Source LLMs and 273 from proprietary LLMs (gpt35-turbo-4k, gpt4o-8k, and gpt4o). Some failed outputs include copying the assurance case template and malformed structures, such as only containing MainClaim, missing SubClaim or containing no evidence element.
A.3 Multi-hop Inference
FT
All Multi-hop inference models are fine-tuned for 10000 steps, 32 batches, 1e-05 learning rate, and weight decays 0.01. The models are optimised with a 1e-05 learning rate, and the weights decay by 0.01.